OpenStudy (anonymous):

What is the volume of a pyramid with base area B = 6 ft2 and height h = 13 ft?

OpenStudy (sidsiddhartha):

\[volume of a pyramid=(1/3)*b ^{2}*h\] your b^2=6 and h=13 now just solve it

OpenStudy (anonymous):

156?

OpenStudy (sidsiddhartha):

its 26 vol=(1/3)*6*13=26

OpenStudy (anonymous):

Isnt the base squared tho?

OpenStudy (anonymous):

@sidsiddhartha

OpenStudy (sidsiddhartha):

base area is given=6 area itselt is squared so b=sqrt(6) b^2=6

OpenStudy (anonymous):

Ohh

OpenStudy (sidsiddhartha):

:)

OpenStudy (anonymous):

Will you help me with a few more?

OpenStudy (anonymous):

The base of a square pyramid has a side length of 6 inches. The height of the pyramid is 12 inches. What is the volume of the pyramid? Um, Im getting 72...

OpenStudy (sidsiddhartha):

side is given 6 inches so b=6 and b^2=36 vol=(1/3)*36*12=144

OpenStudy (anonymous):

Ohh ok

OpenStudy (sidsiddhartha):

:)

OpenStudy (anonymous):

One more... Cause, like, I cant get it...XD http://lv.k12.com/mediaFiles/Ninth_Grade/Math/Unit17866/Lesson69397/05_07_LQ_05.gif For this one I do 18*24=? And then that times height and 1/4 right?

OpenStudy (anonymous):

1/3*